Given the following mixture of two compounds 40.00 g of X (MW =92.00 g/mol)(density 0.824 g/mL) and 825.00 mL of Y (69.00 g/mol))(density 0.904 g/mL), calculate the molality of X.
1
Expert's answer
2020-08-31T07:05:36-0400

Here solvent is Y and solute is X

Mass of solvent = 825× 0.904 gm

= 745.8 gm

Moles of X present = 40/92

= 0.435 mole

In 745.8 gm solvent solute present is 0.435 moles

In 1000 gm of solvent, solute present is = (0.435/745.8)×1000

=0.58

Molality of the solution is 0.58 m